OpenStudy (jessiegonzales):

MEDAL!!! ALGEBRA 2 4. Write the equation in slope-intercept form. What are the slope and y-intercept? -9x+10y=-9

OpenStudy (jessiegonzales):

A. y=9/10x+9/10; slope:9/10;y-intercept:9/10 B. Y=9/10X-9/10; slope 9/10;y-intercept:9/10 C. y=-9/10x+9/10;slope 9/10;y-intercept:9/10 D.y=0/10x-9/10;slope9/10;y-intercept;-9/10

OpenStudy (anonymous):

the general equation for a straight line is y=mx+c where m is the slope and c is the y intercept so dividing the whole equation by 10..we get -9x/10+y=-9/10 => y=9x/10-9/10 so the slope is 9/10 intercept is -9/10

OpenStudy (jessiegonzales):

thankyou so much <3